### v3

# PixT3: Pixel-based Table-To-Text Generation

Abstract
Table-to-text generation involves generating appropriate textual descriptions given structured tabular data. It has attracted increasing attention in recent years thanks to the popularity of neural network models and the availability of large-scale datasets. A common feature across existing methods is their treatment of the input as a string, i.e., by employing linearization techniques that do not always preserve information in the table, are verbose, and lack space efficiency. We propose to rethink data-to-text generation as a visual recognition task, removing the need for rendering the input in a string format. We present PixT3, a multimodal table-to-text model that overcomes the challenges of linearization and input size limitations encountered by existing models. PixT3 is trained with a new self-supervised learning objective to reinforce table structure awareness and is applicable to open-ended and controlled generation settings. Experiments on the ToTTo Parikh et al. (2020a) and Logic2Text Chen et al. (2020c) benchmarks show that PixT3 is competitive and, in some settings, superior to generators that operate solely on text. [^1] Our code, models, and data are available at https://github.com/alonsoapp/PixT3.
PixT3: Pixel-based Table-To-Text Generation
Iñigo Alonso, HiTZ Center - Ixa, University of the Basque Country UPV/EHU inigoborja.alonso@ehu.eus                        Eneko Agirre HiTZ Center - Ixa, University of the Basque Country UPV/EHU e.agirre@ehu.eus                        Mirella Lapata Institute for Language, Cognition and Computation, University of Edinburgh mlap@inf.ed.ac.uk
1 Introduction
Generating text from structured inputs such as tables, tuples, or graphs, is commonly referred to as data-to-text generation (Reiter and Dale, 1997; Covington, 2001; Gatt and Krahmer, 2018). This umbrella term includes several tasks ranging from generating sport summaries based on boxscore statistics (Wiseman et al., 2017), to producing fun facts from superlative Wikipedia tables Korn et al. (2019), and creating textual descriptions given biographical data (Lebret et al., 2016). From a modeling perspective, data-to-text generation is challenging as it is not immediately obvious how to best describe the given input. For instance, the table in Figure 1 can be verbalized in different ways, depending on the specific content we choose to focus on. In controlled data-to-text generation Parikh et al. (2020a), models are expected to generate descriptions for pre-selected parts of the input (see the highlighted cells in Figure 1).
Regardless of the generation setting, numerous approaches have emerged in recent years with different characteristics. A few exploit the structural information of the input (Puduppully et al., 2019; Chen et al., 2020b; Wang et al., 2022), use neural templates (Wiseman et al., 2018), or resort to content planning (Su et al., 2021; Puduppully et al., 2022). While others (Chen et al., 2020a, c; Aghajanyan et al., 2022; Kasner and Dusek, 2022) improve on fluency and generalization by leveraging large-scale pre-trained language models (Devlin et al., 2019; Raffel et al., 2020). A common feature across these methods is their treatment of tabular input as a string, following various linearization methods. As an example, Figure 1 shows the representation of tabular data (top) as a sequence of (Column, Row, Value) tuples (bottom).
Problematically, representing tabular information as a linear sequence results in a verbose representation that often exceeds the context window limit of popular Transformer models Vaswani et al. (2017). The challenge of processing such long sequences has fostered the development of even more controlled methods which refrain from encoding the table as a whole, concentrating exclusively on highlighted content (e.g., only the yellow cells in Figure 1). Unfortunately, models trained on abridged input have difficulty generalizing to new domains while being practically ineffective in scenarios where content selection is not provided.
Table Title: Shuttle America
Section Title: Fleet
[…]
Target Description: Shuttle America operated the E-170 and the larger E-175 aircraft for Delta Air Lines.
Figure 1: Example of table-to-text generation taken from the ToTTo dataset Parikh et al. (2020a). In the controlled setting, a natural language description is generated only for highlighted (yellow) cells. The table is linearized by encoding each value as a (Column, Row, Value) tuple. We only show the first row, for the sake of brevity.
In this paper we propose to rethink data-to-text generation as a visual recognition task, allowing us to represent and preserve tabular information compactly. Vision Transformers (ViTs; Dosovitskiy et al. 2021) have significantly advanced the field of visual language understanding Kim et al. (2022); Davis et al. (2022) demonstrating proficiency in various tasks, including language modeling Rust et al. (2023), visual document understanding Huang et al. (2022), and visual question answering Masry et al. (2022). Our work builds on Pix2Struct Lee et al. (2023), a pretrained image-to-text model which can be fine-tuned for visually-situated language tasks. We recast data-to-text generation as an image-to-text problem and present PixT3, a Pixel-based Table-to-Text model, which is generally applicable to open-ended and controlled generation settings, overcoming the challenges of linearization and input size limitations encountered by existing models.
Our contributions can be summarized as follows: (a) we introduce the first pixel-based model for table-to-text generation and showcase its robustness across generation settings with varying table sizes; (b) we propose a new training curriculum and self-supervised learning objective to reinforce table structure awareness; (c) automatic and human evaluation results on the ToTTo benchmark Parikh et al. (2020b) show that PixT3 excels in open-ended generation, leading to improved faithfulness and generation quality, while being competitive with existing methods in controlled scenarios; and (d) we present a new dataset based on Logic2Text Chen et al. (2020c), which allows us to evaluate generalization capabilities of current table-to-text models.
2 Related Work
The bulk of previous work treats tables as textual objects. Several techniques have been developed to extract accurate information from them (Puduppully et al., 2019; Chen et al., 2020b) using templates (Wiseman et al., 2018), enforcing table structure awareness Mahapatra and Garain (2021); Wang et al. (2022), applying contrastive learning An et al. (2022); Chen et al. (2023b) or focusing on content planning (Su et al., 2021; Puduppully et al., 2022). Other techniques (Chen et al., 2020a, c; Aghajanyan et al., 2022; Kasner and Dusek, 2022) improve fluency and generalization by leveraging large-scale pretrained language models (Devlin et al., 2019; Raffel et al., 2020). Tables are generally linearized, even when special-purpose techniques are developed for encoding table structure Wang et al. (2022). Dedicated table understanding techniques Wang et al. (2021); Jin et al. (2023) eschew linearization but have not been integrated with generation tasks.
Previous attempts to address table-to-text generation from a visual recognition perspective Dash et al. (2023); Srihari et al. (2003) have relied on OCR methods which first extract text from the image and then feed it as a string to a generation model. Aside from being noisy, these techniques typically embrace a text-centric point of view, treating the image as a limitation rather than an informative modality. Our work builds on recent visual language understanding models Kim et al. (2022); Davis et al. (2022); Lee et al. (2023) which are based exclusively on pixels and have managed to outperform OCR methods in several natural language processing tasks Rust et al. (2023); Huang et al. (2022); Masry et al. (2022); Salesky et al. (2023).
The field of Vision Language Models (VLMs) has also experienced significant growth in recent years Liu et al. (2023); Ye et al. (2023b); Bai et al. (2023); Wang et al. (2023); Alayrac et al. (2022). While most of them focus primarily on natural images, a few are starting to explore the application of dual encoder architectures to visually represented language Ye et al. (2023a); Zhang et al. (2023). However, these architectures are not parameter lean (with increased model size of a factor of 40 or more compared to Pix2Struct), and some continue to rely on fixed resolution images which can be particularly problematic when processing tabular data.
A few other efforts have recently explored multimodal approaches to processing tables for various tasks, including table-to-text generation. Dash et al. (2023) convert images into HTML tokens which are subsequently linearized and processed by a traditional text-to-text model. Other work Chen et al. (2023a) focuses on recognizing the structure of tables from images as an independent task. It also leverages multimodal pretraining and unsupervised table structure learning objectives, but ignores the content of table cells and their relations. To the best of our knowledge, our work is the first to conceptualize data-to-text generation as a visually-situated language understanding problem.
3 Problem Formulation
The task of table-to-text generation aims to take a structured table $\bm{t}$ as input and output a natural language description $\bm{y}=[y_{1},\dots,y_{k}]$ where $k$ is the length of the description. Table $\bm{t}$ is typically reformatted as a sequence of textual records $\bm{t}=[t_{1,1},t_{1,2},\dots,t_{i,j},\dots,t_{m,n}]$ where $m$ and $n$ respectively denote the number of rows and columns of $\bm{t}$.
We approach this task from a visual recognition perspective, and expect the input table to be an image $\bm{x}$. The image is reshaped into a sequence of patches analogous to linguistic tokens. More formally, for an input image $\bm{x}\in R^{H\times W\times C}$ and patch size $p$, we create $N$ image patches denoted as $x_{p}\in R^{N\times(P^{2}\cdot C)}$. $(H,W)$ is the resolution of the original image, $C$ is the number of channels, $(P,P)$ is the resolution of each image patch, and $N=\frac{HW}{P^{2}}$ the resulting number of patches, which serves effectively as the input sequence length. Our proposed model learns to autoregressively estimate the conditional probability of a text sequence from a source image as:
$P(\bm{y}|\bm{x};\bm{\theta})=\prod\limits_{i=1}^{n}P(y_{i}|\bm{y}_{<i},\bm{x};\bm{\theta})$ (1)
where ${\bm{\theta}}$ are transformer parameters and $\bm{y}_{<i}$ the words decoded thus far.
We further define three generation settings, which manipulate the information provided to the model in terms of content selection (see Appendix B for visualization). In the tightly-controlled setting (TControl), the model is given highlighted cells only, ignoring the table. Most recent approaches benchmark model performance in this setting Wang et al. (2022); An et al. (2022); Chen et al. (2023b); Su et al. (2021); Kale and Rastogi (2020). In the loosely controlled setting (LControl), the model is given highlighted cells and the entire table. This is the original setting for which the ToTTo dataset Parikh et al. (2020a) was constructed. Finally, we introduce the open-ended setting (OpenE), where the model is given the table without any highlighting.
Figure 2: Overview of PixT3 generation model.
4 The PixT3 Model
PixT3 is an image-encoder-text-decoder model based on Pix2Struct Lee et al. (2023). It expects image rendered tables and generates descriptions thereof (see Figure 2). Pix2Struct is a Vision Transformer model pretrained on 80M screenshots of web pages extracted from URLs in the C4 corpus Raffel et al. (2020). It splits input images into patches of 16$\times$16 pixels (see Figure 2), linearly embeds each patch, adds position embeddings, and feeds the resulting sequence of vectors to a standard Transformer encoder Vaswani et al. (2017).
Pix2Struct was first warmed up with a reading curriculum Rust et al. (2023); Davis et al. (2022), to improve training stability and fine-tuning performance and then pretrained with a screenshot parsing objective; specifically, it generates a simplified version of an HTML subtree that represents a highlighted area of a web page screenshot. It also adds a BART-like Lewis et al. (2020) learning signal to pretraining by masking 50% of the text in the input and then requiring the model to produce the entire subtree. Importantly for our table-to-text generation task, Pix2Struct supports variable image resolution and multiple aspect ratios. It first re-scales the input (up or down) to extract the maximal number of fixed-size patches that fit within a given sequence length and then replaces the typical 1-dimensional absolute positional embedding with a 2-dimensional one, which adds resolution flexibility and removes any aspect ratio distortion.
We initialize PixT3’s model weights with Pix2Struct; we next adopt a curriculum training strategy which instills in our model knowledge about tables and their structure (see Section 4.2); and finally, we fine-tune on table-to-text generation datasets such as ToTTo Parikh et al. (2020a) with a task-specific supervised objective.
4.1 Table-to-Image Rendering
We parse tables to HTML, and subsequently render them into images. We also render table metadata (e.g., Wikipedia page and section title), if it exists, as part of the image, adding it on top of the table. Tables are rendered into three different images corresponding to the generation settings defined in Section 3 (see Appendix B, Figure 6).
Although Pix2Struct can handle variable resolutions and input patches, very long inputs are nevertheless computationally expensive. Following Lee et al. (2023), we set the maximum input length to 2,048 patches (of 16$\times$16 pixels) which corresponds to a maximum image size of 524,288 pixels. 41.74% of the tables in a dataset like ToTTo Parikh et al. (2020a) exceed this size (see Figure 5 in Appendix A), with 5% being larger than 8.3M pixels (32,768 patches). Indiscriminately down-scaling all images exceeding the maximum input length would negatively affect performance, especially for very big tables, effectively rendering them unreadable (we showcase how image size affects model performance in Figure 4). To avoid this as much as possible, we truncate the image to fit within a maximum down-scaling factor $\gamma$. In other words, images are first compressed to $\gamma\%$ of their original size and then truncated from left to right until they fit into 2,048 patches. The optimal value for $\gamma$ is determined empirically (see Appendix C).
4.2 Structure Learning Curriculum
Pix2Struct is a general-purpose visual language understanding model, and as such it is not particularly knowledgeable about tables and their structure. Tables can be presented in a variety of ways visually, such as spanning multiple columns or rows, with or without horizontal and vertical lines, non-standard spacing and alignment, and text formatting. Aside from presentation, there are various conventions about the underlying semantics of tables and their structure, e.g., each cell is only related to cells in the same column and row. These challenges have led to the development of dedicated table understanding techniques Jin et al. (2023); Wang et al. (2022) in the domain of text but cannot be readily ported to images.
Instead, we encourage PixT3 to adhere to tabular conventions, by first training it on an intermediate supporting task. This acts as a structure learning curriculum, exposing the model to the rules governing tables. We next elaborate on the intermediate task, its corresponding dataset, and the proposed self-supervised objective.
Dataset for Intermediate Training
Existing datasets like ICDAR2021 Kayal et al. (2021) and TableBank Li et al. (2019) are representative of the task of parsing table images into their structure and, in theory, could be used for our intermediate training purposes. However, they focus on scientific tables which do not follow the typical distribution of Wikipedia tables found in ToTTo Parikh et al. (2020a), e.g., in terms of size and cells spanning across rows and columns. We instead propose to create a synthetic image-to-text dataset, making use of the table rendering pipeline described in Section 4.1. Although we generate tables specifically tailored for our use-case, the generation process is flexible and can be adapted to other domains with different characteristics.
We determine the structure of each table (size, column, and row spans) randomly, following ToTTo’s training set distribution. We cap the generation process at a maximum of 20 columns and 75 rows. Table cells are filled with synthetic values consisting of a random combination of one to five random English alphabet characters and digits, functioning as identifiers rather than meaningful values (see Figure 3 for an example). Our dataset contains 135,400 synthetic tables, 120,000 for training, 7,700 for validation, and 7,700 for testing.
Figure 3: Synthetically generated table with a highlighted cell and corresponding pseudo-HTML target sequence (for self-supervised objective). Cells within the target sequence are highlighted in the table with a colored background. For details on the structure of the target, please refer to Appendix D.
Self-supervised Objective
While masking is a widely adopted learning objective Devlin et al. (2019), it does not naturally transfer to our table-to-text generation task; table values are not naturally correlated to neighboring values and thus a masked cell cannot be easily predicted from other cells in its context. Table values could be rearranged so that they correlate to their neighbors, however, early experiments showed that this type of objective does not improve downstream task performance (see Appendix D for details). Another common pretraining objective is table linearization Chen et al. (2023a), which, however, scales poorly with table size, leading to slow pretraining.
We propose a self-supervised objective that encourages PixT3 to capture the relations between cells within a table while generating a small amount of tokens. Specifically, we highlight a random cell in a synthetically generated table, and train the model to produce a sorted list of cells within the same column and row (see Figure 3). Our objective encapsulates a loose notion of table structure, nudging the model to pay attention to the arrangement of columns and rows around a cell. We follow the same pseudo HTML notation introduced in Pix2Struct to format our output sequence, easing the model’s transition from its original screenshot parsing objective to this new one. Note that we consider tables with a heterogeneous structure where cells can span across multiple columns and rows. In such cases, the expected sequence will contain all cells in related rows and columns surrounding the highlighted cell (see Figure 3).
4.3 PixT3 Fine-tuning
The intermediately pre-trained PixT3 is subsequently fine-tuned on an image-rendered dataset (see Section 4.1). In experiments, we use ToTTo Parikh et al. (2020b), however, our approach is not tied to a particular style of tables. Due to our model’s requirement for unimodal input, we treat table-related information (such as its title) as part of the table itself and render them both as one image (see Lee et al. 2023 for a similar approach).
5 Experimental Setup
Model Configuration
All our experiments were conducted with the base pretrained Pix2Struct [^2] https://github.com/google-research/pix2struct model (282M parameters). We trained PixT3 variants for the three table-to-text generation settings defined in Section 3. All PixT3 models were fine-tuned on ToTTo Parikh et al. (2020a) with tables rendered as images following the procedure outlined in Section 4.1. The maximum down-scaling factor $\gamma$ was set to 0.39.
PixT3 models were fine-tuned with a batch size of 8 and a gradient accumulation of 32 steps on a single NVIDIA A100 80GB GPU. Checkpoints were selected according to best performance on the validation set. All models used an input sequence length of 2,048 patches and were optimized with AdamW Loshchilov and Hutter (2017). We used a learning rate scheduler with a linear warmup of 1,000 steps to 0.0001, followed by cosine decay to 0. The decoder maximum sequence length was set to 50 tokens, which covers 97.49% of the target descriptions in the training data. PixT3 was trained for 1.4k steps with the self-supervised objective described in Section 4.2. Our decoder was not frozen during intermediate training, as initial experiments showed that a fully trained model outperformed one with frozen decoder weights. A full list of fine-tuning hyper-parameters can be found in Appendix H.
Datasets
We evaluated our model on ToTTo Parikh et al. (2020a), a large-scale, manually curated dataset representative of several domains and types of tables. We also assessed the generalization capabilities of PixT3 on out-of-distribution tables. We created an out-of-domain benchmark with content selection annotations similar to ToTTo based on Logic2Text Chen et al. (2020c), an existing dataset which contains a total of 10,161 Wikipedia tables, paired with human-authored descriptions and logical forms. Logic2Text differs from ToTTo in that descriptions are not simple verbalisations of table rows and columns, but require some form of reasoning (e.g., comparisons or counting operations). We were able to automatically trace values mentioned in the logical form back to the cells of the input tables Alonso and Agirre (2023), thus obtaining highlighted cell annotations similar ToTTo’s (see Appendix E for an example). We report results on the official test set (1,085 examples).
Model Comparison
We evaluated PixT3 against several text-only models with similar parameter sizes. These include CoNT An et al. (2022), the top performant (published) model in the ToTTo leaderboard. [^3] A model named SKY appears to slightly outperform CoNT in the leaderboard, however, at the time of writing, we were not able to verify this, i.e., by finding a publication or preprint describing this model. CoNT is a text-to-text generation model which makes use of contrastive learning, through improved selection of contrastive examples, a new contrastive loss, and a global decoding strategy. CoNT expects the input table to be converted to a string, and is built on top of T5-base (220M parameters). We also compared against Lattice Wang et al. (2022), a model which enforces awareness of table layout though pruning the attention flow and encoding cells in a way that is invariant to their relative position in a sequence. This model also uses T5-base and expects linearized input. In addition, we report results with vanilla T5-base which performed competitively on the ToTTo leaderboard without any task specific modifications Kale and Rastogi (2020); An et al. (2022). All comparison models and PixT3, were trained on the ToTTo training set in our three generation settings. [^4] Comparison models were trained with the authors’ publicly available scripts.
For our out-of-domain experiments, we also compare against LLaVA-1.5 Liu et al. (2023), a large pretrained multimodal model (13B parameters) which is built on top of the CLIP visual encoder Radford et al. (2021) and the Vicuna-7B language model Zheng et al. (2023), and fine-tuned on vision-language instructions. LLaVA has not been fine-tuned specifically for table-to-text generation, however, it is interesting to see if sufficiently large scale is all it takes to do well on the table-to-text generation task. LLaVA can only handle a single image at each forward pass. This limitation prevents it from performing inference in an in-context learning setting, where the model has access to multiple input-output examples at the same time. To approximate in-context learning as closely as possible, we provided LLaVA with an image, an instruction, and three table descriptions as output examples for each generation setting (see Appendix F for details). We summarize the number of parameters for all comparison models in Table 2.
we do still provide a few description examples in our prompt to ensure a fair zero-shot comparison. All prompts used for LLaVA in this evaluation can be found in Appendix F.
Dev TestN TestO Test
Model BL PR BL PR BL PR BL PR
T5-base 47.7 57.1 38.9 51.2 55.4 61.1 47.2 56.2
T5-3B 48.4 57.8 39.3 51.6 55.1 60.7 47.2 56.2
Lattice 48.0 58.4 40.0 53.8 55.9 62.4 48.0 58.1
CoNT 49.0 58.6 40.6 53.7 56.7 62.5 48.7 58.1
TControl PixT3 45.7 55.7 37.5 50.6 53.2 60.4 45.4 55.5
T5-base 24.5 27.2 19.4 23.9 29.4 30.3 24.5 27.1
T5-3B 23.6 26.0 18.0 22.4 28.7 29.2 23.4 25.8
Lattice 24.9 31.0 20.8 27.7 27.5 33.8 24.4 30.8
CoNT 23.8 29.3 19.2 26.1 28.7 32.3 23.9 29.2
LControl PixT3 46.2 55.1 38.1 50.3 52.7 59.0 45.4 54.7
T5-base 21.5 23.5 16.8 21.0 26.5 26.5 21.7 23.8
T5-3B 20.8 22.9 16.7 20.3 25.5 25.5 21.2 22.9
Lattice 20.9 26.1 17.6 24.3 23.7 27.6 20.8 25.9
CoNT 21.7 25.8 16.9 23.2 26.3 28.3 21.6 25.8
OpenE PixT3 24.8 28.3 20.5 26.3 28.9 30.3 24.7 28.3
Table 1: Automatic evaluation results on ToTTo in three generation settings: tightly controlled (TControl), loosely controlled (LControl), and open-ended (OpenE). We report BLEU (BL) and PARENT (PR) results on the development (Dev) and Test sets, including the overlapping (TestO) and non-overlapping (TestN) test set splits. BLEURT results are in Appendix E.
6 Results
PixT3 is the best performing model in loosely controlled and open-ended generation settings.
Table 1 summarizes our results on ToTTo in our three generation settings. We evaluated model performance automatically with the same metrics used to rank participant systems in the ToTTo leaderboard. These include BLEU Papineni et al. (2002) which is as a proxy for fluency, PARENT Dhingra et al. (2019), a metric proposed specifically for data-to-text evaluation that takes the table into account, serving as a proxy of faithfulness, and BLEURT Sellam et al. (2020); the latter is a composite metric that takes a reference and model output as input, and returns a score that indicates the extent to which the output is fluent and conveys the meaning of the reference. Note that ToTTo features two splits in the development/test set containing tables whose header values are present (overlapping split) and absent (non-overlapping split) in the training set. Results on the test set, which is not publicly available, were obtained via submitting to the ToTTo leaderboard.
We first discuss our results on the tightly controlled generation setting (TControl) where models are not given the full table, just the highlighted cells. We would not expect PixT3 to excel at this setting, which is better suited to text-to-text models (highlighted cells make for non-descriptive images, see Appendix B, Figure 6). PixT3 is indeed unable to outperform CoNT, Lattice, and related T5 variants, falling 3.5 BLEU points behind on the development set and 3.7 on the test set. However, LControl, the loosely controlled generation setting, better showcases the advantages of PixT3, which in this case demonstrates almost a two times improvement over CoNT and T5 models. Performance degrades drastically for all systems in the open-ended setting (OpenE) which is challenging; models are expected to perform content selection in addition to text generation, and could produce table descriptions which are valid but different from the reference. Automatic metrics based on n-gram overlap are particularly punitive in this case. Nevertheless, PixT3 is superior to CoNT, Lattice, and T5 across evaluation metrics.
Model Size BLEU PARENT
LLaVA 13B 12.6 34.36
T5-base 220M 16.8 55.97
T5-3B 3B 17.7 52.75
Lattice 220M 19.8 61.05
CoNT 220M 18.8 61.73
TControl PixT3 282M 20.6 61.86
LLaVA 13B 5.9 23.18
T5-base 220M 11.5 40.02
T5-3B 3B 10.9 35.45
Lattice 220M 11.5 40.02
CoNT 220M 11.8 43.25
LControl PixT3 282M 21.5 56.45
LLaVA 13B 6.7 20.14
T5-base 220M 7.9 30.67
T5-3B 3B 9.5 29.47
Lattice 220M 11.7 38.12
CoNT 220M 11.0 36.94
OpenE PixT3 282M 11.4 35.68
Table 2: Automatic evaluation results on Logic2Text in three generation settings: tightly controlled (LControl), loosely controlled (LControl), and open-ended (OpenE). All models (except LLAVA) were fine-tuned on ToTTo and tested on Logic2Text. BLEURT results are in Appendix E.
PixT3 generalizes to out-of-domain tables which require reasoning skills.
We next evaluate whether PixT3 generalizes to unseen tables, outside ToTTo’s distribution. Table 2 shows our results on Logic2Text Chen et al. (2020c), again following the three generation settings. Compared to ToTTo, Logic2Text is a more challenging dataset as most descriptions rely on reasoning over the entire table. This results in poor model performance in the TControl setting which does not include the table as input. Nonetheless, we observe that PixT3 excels at the LControl setting, even though it has to process and reason over the entire table. The OpenE setting is challenging for all models as they are asked to identify interesting cells to talk about in out-of-domain tables. PixT3 still maintains an edge over T5 and LLaVA, performing on par with CoNT and Lattice. We observe that LLaVA cannot match the performance of PixT3 and T5-based models. This underscores the importance of task-specific fine-tuning over parameter size. We present output examples in Appendix E.
Figure 4: Model performance (CoNT, T5, PixT3, Lattice, and PixT3 with 512 patch input size) in the LControl setting across 18 table size groups (logarithmic scale). Upper and lower bounds in shaded areas correspond to results for the overlapping and non-overlapping ToTTo splits, while central points correspond to results overall. We report results with PARENT, other metrics show similar tendencies. We refer to Appendix A for further details.
PixT3 is robust against table input size.
In Figure 4, we analyze the effect of table size on model performance. As can be seen, T5, Lattice, and CoNT are severely affected: the bigger the table, the less accurate the generated description. PixT3 is evidently more robust, showing degradation in performance only for very big tables. We also examined whether PixT3 has an edge because of its ability to encode longer inputs. Recall that CoNT, Lattice, and T5-base utilize a fixed input length of 512 tokens, while PixT3 uses 2,048 patches. We thus trained a PixT3 variant with input length set to 512 patches. As shown in Figure 4, the more constrained PixT3 model is slightly worse and more likely to degrade with increased table size but consistently outperforms CoNT, Lattice, and T5.
The structure learning curriculum improves generation quality across metrics.
In Table 3 we perform an ablation study comparing PixT3 with and without our structure learning curriculum and self-supervised objective (Section 4.2). For both models we follow the same fine-tuning process: we render tables into images, identify the optimal point of image compression and truncation (see Section 4.1), and perform hyper-parameter search to optimize Pix2Struct-base for our task. Vanilla PixT3 (second row in Table 3) shows a substantial improvement over an out-of-the-box Pix2Struct model which achieves a BLEU score of 0.2 and PARENT score of 0.6 on the ToTTo development set. Adding the intermediate training curriculum (second row in Table 3) slightly improves vanilla PixT3 across evaluation metrics.
Manual inspection of the descriptions produced by the two PixT3 model variants reveals they are often semantically equivalent to the target (43% of the time). Nevertheless, the intermediate training curriculum substantially reduces structure-based faithfulness errors, especially in the OpenE setting. On a sample of 200 outputs (randomly selected from the development set), we found that 23% of the descriptions produced by vanilla PixT3 disregard or misinterpret the structure of the table. Structural faithfulness errors reduce to 7% when PixT3 is trained with our structure learning curriculum.
Dev Test
Models BL PR BRT BL PR BRT
Pix2Struct 0.2 0.6 --1.433
PixT3 (W/o SLC) 38.7 46.0 --0.003 38.3 45.6 0.001
PixT3 (With SLC) 39.2 46.5 0.008 38.7 46.3 0.007
Table 3: PixT3 with and without structure learning curriculum (SLC); we report results on the ToTTo development (Dev) and Test set with BLEU (BL), PARENT (PR), and BLEURT (BRT), averaged across the three generation settings.
PixT3 is most faithful in loosely controlled and open-ended generation settings.
We further conducted a human evaluation study to quantify the extent to which the generated descriptions are faithful to the table. We evaluated PixT3, and the two best performing text-only systems (CoNT, and Lattice) on two sets of 100 randomly selected table-description pairs from ToTTo (development set) and Logic2Text (test set), in the three generation settings. Crowdworkers were presented with an uncompressed image of a table, its page and section title, and a model generated description. As an upper bound, we also elicited judgments for the human curated reference descriptions for the same ToTTo and Logic2Text examples. Participants were asked to determine whether a description was "True" or "False" based on the information provided in the table and/or its title and subtitle (see instructions in Appendix G). Overall we elicited 7,200 judgments (100 examples $\times$ 3 generation settings $\times$ 4 model descriptions $\times$ 3 participants $\times$ 2 datasets). Crowdworkers were recruited using the online platform Prolific. [^5] https://www.prolific.com
Table 4 shows the results of the human evaluation, specifically the proportion of descriptions deemed faithful. As expected, the human authored Reference description is consistently faithful across generation settings. CoNT is more faithful in TControl but deteriorates in the LControl and OpenE settings. We further examined whether differences among systems are statistically significant using paired bootstrap resampling. PixT3 is significantly worse ($p<0.05$) than the Reference in TControl but not CoNT or Lattice. In LControl all differences between systems are statistically significant ($p<0.05$). In OpenE, PixT3 is significantly different ($p<0.05$) from CoNT and Lattice but not from the Reference. Inter-rater agreement was moderate with a Fleiss’ Kappa coefficient of 0.55 (Fleiss, 1971).
Model TControl LControl OpenE
Reference 87 84 89
Lattice 79 16 20
CoNT 76 16 35
ToTTo PixT3 69 72 78
Reference 81 87 86
Lattice 34   3 16
CoNT 35   3 26
L2T PixT3 32 40 60
Table 4: Human evaluation results on ToTTo and Logic2Text (L2T). Proportion of descriptions rated as faithful for PixT3, CoNT, and Reference in three generation settings: tightly controlled (LControl), loosely controlled (LControl), and open-ended (OpenE).
7 Conclusion
In this paper, we leverage the capabilities of Vision Transformers to recast table-to-text generation as a visual recognition task, removing the need for rendering the input in a string format. Our model, PixT3, introduces a new training curriculum and self-supervised learning objective in order to capture the structure and semantics of tables. Experiments across constrained and open-ended generation settings show it is robust to different table sizes, performing competitively and often better than state-of-the-art models. PixT3 is also able to handle new domains with unseen tables, as evidenced by our results on Logic2Text, a new dataset which we propose for assessing the generalization capabilities of table-to-text generation models.
Avenues for future research are many and varied. There are several downstream tasks which stand to benefit from a pixel-based view of textual information, including multilingual table-to-text generation, and semantic parsing. We would also like to investigate additional objectives and inductive biases that can better capture the structure of tables and inter-cell dependencies.
8 Limitations
While PixT3 shows promising results, its performance is affected by the dimension of the input tables (for instance, 16% of the Wikipedia tables in ToTTo remain too big for PixT3 to represent effectively). It would be interesting to look into alternative ways of preprocessing very large tables, e.g., by rendering them via multiple images. While our proposed intermediate training methodology mitigates faithfulness errors, the model still struggles with hallucinations, falling short of human-level performance.
Finally, PixT3, as well as other comparison systems, have limited reasoning capabilities, e.g., they cannot infer information which is not explicitly stated in the table or make logical connections between concepts. PixT3’s superior performance in terms of faithfulness on Logic2Text (see Table 4) is due to generating simpler sentences rather than superior reasoning skills. Thus, aside from new training objectives, a promising direction would be to combine the visual representations with an intermediate planning component that encourages the model to reason about the input while generating the output.
Acknowledgements
We thank the meta-reviewer and anonymous reviewers for their constructive feedback. The authors also thank Ander Salaberria for his insightful comments on earlier versions of this work. We gratefully acknowledge the support of the UK Engineering and Physical Sciences Research Council (grant EP/L016427/1), the Basque Government (Research group funding IT-1805-22), MCIN/AEI/10.13039/501100011033 project AWARE (TED2021-131617B-I00), European Union NextGenerationEU/PRTR, and the LUMINOUS project (HORIZON-CL4-2023-HUMAN-01-21-101135724).
Appendix A Table Size Distribution in ToTTo
We measure the size of a table by the total amount of pixels in its corresponding rendered image. We then calculate the distribution of each size, and group tables into 20 buckets accordingly. Each bucket covers a logarithmically increasing amount of table sizes. Figure 5 shows the resulting buckets and the proportion of ToTTo examples in each (development set). The quality of descriptions generated within each group, are evaluated in Section 6, see Figure 4.
Figure 5: Proportion of ToTTo examples (development set) per table size (shown in logarithmic scale).
Appendix B Table-to-Text Generation Settings
Figure 6 illustrates how the image input to PixT3 differs according to three generation settings: tightly controlled (the model is given only highlighted cells, no table), loosely controlled (the model is given the table and highlighted cells), and open-ended (the model is given the table without any highlighting).
Figure 6: PixT3 input image examples (and reference) in three generation settings: tightly controlled (TControl), loosely controlled (LControl), and open-ended (OpenE).
Appendix C Image Truncation and Down-scaling
We explored the impact of down-scaling on model performance and its tradeoff with truncation. We conducted a series of experiments wherein PixT3 models were trained on versions of ToTTo with varying down-scaling factor $\gamma$: 0.87, 0.58, 0.39, 0.26, and 0.00. Note that $\gamma$=0.00 corresponds to a setting where no truncation takes place, only down-scaling. According to the results shown in Table 5, it is best to combine truncation with down-scaling, none of the extreme settings (no truncation vs too much truncation) are beneficial. The optimal $\gamma$ value is 0.39.
Epoch $\gamma$ 0.00 0.26 0.39 0.57 0.87
16 28.71 29.13 29.47 29.58 27.47
17 28.99 29.53 29.99 29.70 27.69
18 29.67 30.04 30.55 30.21 28.13
19 29.98 30.04 30.63 30.54 28.33
20 29.83 30.21 30.68 30.53 29.39
Table 5: Evaluation results (BLUE) for PixT3 model in tightly controlled generation setting for different $\gamma$ down-scaling factors. We show the Last five epochs on the ToTTo training set.
Appendix D Intermediate Training
Synthetic Dataset Generation
In this section we provide a more detailed description regarding the generation of synthetic tables for intermediate training. As our goal was to generate tables with a structure similar to ToTTo, we first measured the probability distribution of columns, rows, column spans and row spans for the tables in the training set to avoid over-fitting and contamination. We observed that the distribution of columns (up to 20 columns) remained almost constant across tables, and did not affect the probability distribution of rows. As a result, we aggregated row numbers across columns and computed a single distribution for rows to simplify our generation task, using discrete probability distributions. In order to limit the size of the generated tables we cap the number of columns and rows to 20 and 75, respectively. For the synthetic text within the cells, we randomly generated digits in the [1–5] range and character sequences from [A–Z, a–z] which gave us a total of 776,520,240 permutations of possible unique cell values.
Overall, we generated 120K tables accompanied with target pseudo HTML descriptions. The latter were on average 121 tokens long, with the longest sequences containing 877 tokens. In experiments, we observed that text size affects mainly the average count of tokens, whereas the number of table columns and rows influences the length of the target sequences. The sequences follow a hierarchical structure defined by the characters < and >. In the first hierarchical level, one container can be found for each highlighted cell in the table. Each container includes, in the following order, the highlighted cell, the cells in all related columns, and all cells in all related rows. This structure can represent multiple related columns and rows per highlighted cell, as well as multiple highlighted cells per table.
Alternative Objectives
We conducted a set of experiments to identify the best self-supervised objective for our structure learning curriculum. In addition to the objective presented in Section 4.2, we also experimented with a masking objective. Specifically, given a randomly generated table, we filled each cell with text indicative of its position in the table. We then masked random cells and the model was trained to predict the missing cell values (see Figure 7 for an example). We empirically observed that this objective led to worse performance compared to PixT3, even though it resulted in relatively fast training, since the table can be converted into a sequence with a small number of tokens. We hypothesize that this objective only weakly enforces table structure learning as the model does not need to pay attention to all the cells in a column and row to guess the missing value but simply rely on its closest neighbors. We also experimented with a combination of the masking objective discussed here and the structure learning objective described in Section 4.2. However, this model still lagged behind PixT3.
Figure 7: Synthetically generated table with masked cell. Filled cell values denote position in the table.
Dev Set (All) Test Set (Non) Test Set (Over) Test Set (All)
Model BLEURT BLEURT BLEURT BLEURT
T5-base 0.233 0.106 0.354 0.230
T5-3B 0.228 0.104 0.344 0.224
Lattice 0.226 0.103 0.348 0.226
CoNT 0.240 0.116 0.364 0.240
TControl PixT3 0.178 0.044 0.312 0.178
T5-base --0.298 --0.395 --0.191 -0.293
T5-3B --0.309 --0.416 --0.194 -0.305
Lattice --0.287 --0.382 --0.195 -0.288
CoNT --0.293 --0.387 --0.190 -0.289
LControl PixT3 0.169 0.047 0.287 0.167
T5-base --0.371 --0.458 --0.278 -0.368
T5-3B --0.385 --0.456 --0.301 -0.378
Lattice --0.377 --0.451 --0.302 -0.377
CoNT --0.370 --0.452 --0.281 -0.366
OpenE PixT3 --0.332 --0.414 --0.258 --0.336
Table 6: BLEURT results on ToTTo for T5, PixT3, Lattice, and CoNT in three generation settings: tightly controlled (LControl), loosely controlled (LControl), and open-ended (OpenE). In the TControl setting, T5 results are taken from Kale and Rastogi (2020) and CoNT results from An et al. (2022). This table complements results reported in Table 1.
Appendix E Additional Results and Examples
In addition to BLEU and PARENT reported in Tables 1 and 2, we also present results with BLEURT in Table 6 and Table 7. We further show example output on the Logic2Text dataset (zero-shot setting) in Figure 8. In the TControl setting, CoNT struggles to produce a coherent sentence, while PixT3 generates a faithful but not very informative one. This is not surprising as the models receive nothing but the title and highlighted cells, making it extremely difficult to generate the target sentence. In LControl, both models have access to the entire table; however, they still produce a false statement, most likely a consequence of the zero-shot nature of our generation task. Finally, in the less constrained OpenE setting, PixT3 generates a coherent and faithful sentence. While CoNT also produces a fluent sentence, it incurs a faithfulness error when mentioning "(+5)" instead of "(-5)". This is likely due to the performance degradation this model experiences when provided with the full table.
Model BLEURT
LLaVA --1.230
T5-base --1.086
T5-3B --1.079
Lattice --1.060
CoNT --1.103
TControl PixT3 --1.104
LLaVA --1.189
T5-base --1.147
T5-3B --1.167
Lattice --1.147
CoNT --1.159
LControl PixT3 --1.073
LLaVA --1.184
T5-base --1.237
T5-3B --1.196
Lattice --1.231
CoNT --1.231
OpenE PixT3 --1.213
Table 7: Automatic evaluation results on Logic2Text in three generation settings: tightly controlled (LControl), loosely controlled (LControl), and open-ended (OpenE). All models (except LLAVA) were fine-tuned on ToTTo and tested on the Logic2Text. This table complements results reported in Table 2.
Refer to caption Reference: Jim Colbert has the second best number of strokes to par. CoNT (TControl): Jim Colbert led the 1973 U.S. open (golf course) with a score of to par. PixT3 (TControl): Jim Colbert took part in the 1973 U.S. open (golf) tournament. CoNT (LControl): At the 1973 U.S. open (golf), Jim Colbert shot a record of 267 (+1) and finished four strokes ahead of runner-up Lee Janzen. PixT3 (LControl): Jim Colbert had a score of 142. CoNT (OpenE): Gary Player scored 137 (+5) and finished five strokes ahead of runner-up Jim Colbert. PixT3 (OpenE): Gary Player won the 1973 U.S. Open (golf) with a score of 137.
Figure 8: Logic2Text table and model output in three generation settings: tightly controlled (TControl), loosely controlled (LControl), and open-ended (OpenE).
Appendix F LLaVA promts
As mentioned in Section 5, our zero-shot experiments involved comparisons against LLaVA-1.5 Liu et al. (2023), a large pretrained multimodal model (13B parameters). We devised the following prompts for each generation setting:
TControl
"Here are some descriptions based on other highlights of other tables ’chilawathurai had the 2nd lowest population density among main towns in the mannar district .’, ’zhou mi only played in one bwf super series masters finals tournament .’, ’tobey maguire appeared in vanity fair later than mike piazza in 2003 .’. Now write a short description based on the following highlighted cells extracted form a table."
LControl
"Here are some descriptions based on the highlights of other tables not present in the input: ’chilawathurai had the 2nd lowest population density among main towns in the mannar district .’, ’zhou mi only played in one bwf super series masters finals tournament .’, ’tobey maguire appeared in vanity fair later than mike piazza in 2003 .’. Now write a short description based on the highlighted cells in this table following the same style as the example descriptions."
OpenE
"Here are some descriptions from other tables not present in the input: ’chilawathurai had the 2nd lowest population density among main towns in the mannar district .’, ’zhou mi only played in one bwf super series masters finals tournament .’, ’tobey maguire appeared in vanity fair later than mike piazza in 2003 .’. Now write a short description stating something from this table following the same style as the example descriptions."
Appendix G Human Evaluation Guidelines
We provide the full set of instructions presented to crowdworkers for the human evaluation study. Our participants were native English speakers from the United Kingdom and the United States of America, with a 50/50 equal gender split between male and female.
Thank you for taking part in our experiment! You will be presented with a table and a computer-generated description of its content. Your task is to determine whether each description is "True" or "False" based on the information provided in the table and/or its title and subtitle (you will see examples later-on). No expert knowledge is required to perform this task. You should evaluate the descriptions given the information presented in the table, without taking any other information into account (e.g., based on your own knowledge or the web).
Here are some guidelines to help you with your evaluation:
Acronyms: tables often have acronyms which the descriptions might spell out. For example, if the table mentions "TD" and the description correctly spells it out as "touch down," you should not consider this "False" (although the description might be false for other reasons).
Implicit information: the description might mention information that can be inferred but is not explicitly spelled-out in the table. For example, it could mention "steam engines" when the table lists theirs names without explicitly talking about steam engines. In this case, the description should not be considered "False".
- You should evaluate each description independently.
- If the description does not make sense and is impossible to evaluate (usually when summarizing very large tables), you should consider it as "False".
We suggest starting by reading the description and then referring to the table to verify if it aligns with its claims.
This data elicitation study is performed by researchers at [REDACTED]. If you have any questions, feel free to contact [REDACTED]. Participation in this research is voluntary. You have the right to withdraw from the experiment at any time. The collected data will be used for research purposes only. We will not collect any personal information. Your responses will be linked to your anonymous Prolific ID for the exclusive purpose of conducting our experiment.
Appendix H PixT3 Fine-tuning Hyper-parameters
PixT3 models across all three settings (TControl, LControl, OpenE) were fine-tuned using the same hyper-parameters. To prevent over-fitting, we employed early stopping based on the BLEU score computed on the validation set every 250 steps. Table 8 enumerates the specific hyper-parameter values used in PixT3, with all remaining parameters set to the default values defined in Pix2Struct Lee et al. (2023).
Hyperparameter Value
Optimizer AdamW
Learning rate 0.0001
Warm-up steps 1000
Max. input patches 2048
Shuffle train data False
Epochs 30
Train batch size 8
Gradient accum. steps 32
Mixed precision fp16
Evaluation batch size 32
Eval freq. steps 250
Inf. beam search 8 beams
Table 8: Hyperparameters used in PixT3.